### v2

# Model-Based Analysis of Mining Fairness in a Blockchain

Abstract
Mining fairness in blockchain refers to equality between the computational resources invested in mining and the block rewards received. There exists a dilemma wherein increasing the transaction processing capacity of a blockchain compromises mining fairness, which consequently undermines its decentralization. This dilemma remains unresolved even with methods such as the greedy heaviest observed subtree (GHOST) protocol, indicating that mining fairness is an inherent bottleneck in the transaction processing capacity of the blockchain system. However, despite its significance, there have been insufficient research studies that have quantitatively analyzed mining fairness. In this paper, we propose a method for calculating mining fairness. First, we approximated a complex blockchain network using a simple mathematical model, assuming that no more than two blocks are generated per round. Within this model, we quantitatively determined local mining fairness and derived several measures of global mining fairness based on local mining fairness. Subsequently, we validated by blockchain network simulations that our calculation method computes mining fairness in networks much more accurately than existing methods. Finally, we analyzed various networks from the perspective of mining fairness.
1 Introduction
Blockchain is a foundational technology used primarily in decentralized currency systems such as Bitcoin [1]. In blockchain systems, transactions are processed in units known as blocks. The creation of a block involves numerous hash calculations, a process referred to as mining. Nodes that perform mining are called miners. Each miner follows a fork choice rule to identify and extend the main chain. When miners successfully create a block, they may be rewarded via a coinbase transaction, by which they would receive what is known as a block reward. However, these block rewards are obtainable only when the blocks that have been generated become part of the main chain.
Mining fairness refers to equality between the computational resources invested in mining and the resulting block rewards. In other words, it is equality between the proportion of hashrate and the proportion of block rewards (hereafter referred to as the block reward rate). If all blocks were incorporated into the main chain, mining fairness would be achieved because the number of blocks generated by each miner would not be affected by the state of the network. However, in practice, not all blocks are included in the main chain because of blockchain forks, and when blocks are discarded as a result, mining fairness is compromised. Forks can be classified into two types: intentional (malicious) and unintentional. The latter occurs when multiple blocks are generated almost simultaneously. This study addresses mining fairness in the context of unintentional forks.
Figure 1: Schematic of mining fairness, illustrating how it establishes a link between transaction processing capacity (TPS) and decentralization. Even if mining fairness is enhanced by rewarding stale blocks, the nothing-at-stake problem emerges, compromising security.
Mining fairness introduces a tradeoff between the transaction processing capacity and decentralization in blockchain systems (Fig. 1); increasing the transaction processing capacity of a system compromises decentralization. The transaction processing capacity depends on the number of transactions processed per block and the block generation interval. To increase this capacity, one might increase the block size and reduce the block generation interval. However, it is well known that increasing block sizes and reducing block generation intervals result in higher fork rates [2]. As previously observed, an increase in the fork rate undermines mining fairness. If mining fairness is reduced, then some miner groups achieve higher profit rates than others. Consequently, miners with lower profit rates end up leaving the system, whereas those with higher profit rates expand, leading to centralization and reduced decentralization.
The dilemma between transaction processing capacity and decentralization in blockchain systems that arises from mining fairness has not yet been resolved. In this context, mining fairness is an inherent bottleneck in the transaction processing capacity. Here, we demonstrate that the dilemma caused by mining fairness is inherent, using the countermeasures adopted by Ethereum [3] (modified greedy heaviest observed subtree (GHOST) protocol) as an example. Increasing the transaction processing capacity of a blockchain leads to more forks, which, in turn, causes two main problems. First, there is an increased risk of attacks such as 51% attacks and selfish mining [4, 5]. Second, mining fairness is compromised. To address the first problem, Ethereum has introduced the GHOST protocol. On the other hand, to address the second problem, which is regarding the impact on mining fairness, Ethereum partially rewards blocks that cause forks but are not incorporated into the main chain (stale blocks). However, this approach has its own problematic aspects. From another perspective, this implies that even blocks that cause forks can receive block rewards, thereby reinforcing economic incentives for attacks such as 51% attacks and selfish mining. This situation shares the same structure as the nothing-at-stake problem. In fact, it is known that in Ethereum, the risk of attacks that compromise mining fairness, including selfish mining [4], is increased [6, 7, 8]. This indicates that the measures taken by Ethereum do not fundamentally address the problems related to mining fairness.
Thus, it is critical to perform further analyses on mining fairness. One possible approach to these analyses is to perform simulations, which, unfortunately, can be time-consuming and impractical. Consequently, alternative approaches have been explored [9, 10, 11, 12, 13, 14, 15, 16]. However, these methods do not correctly account for how mining fairness is compromised by forks (Section 6.1) and, consequently, lead to analyses that are based on only a weak reflection of real-world systems, making it difficult to derive meaningful solutions from them.
In this study, we propose a model-based calculation method for quantitatively analyzing mining fairness. We approximate a blockchain network using a simplified model, assuming that each round contains at most two blocks. A round $r$ is defined as a unit of time that starts with the creation of a block at height $r$. Essentially, we also assume that at most one fork occurs per round. This modeling approach simplifies the analysis of complex blockchain networks.
Afterward, we validate the accuracy of our proposed model-based calculation in measuring mining fairness by conducting simulation experiments. Owing to computational constraints, validating our method in large-scale networks would be challenging. Instead, we perform the validation in networks with small numbers of miners. Our results demonstrate that the model-based calculation quantitatively determines mining fairness much more accurately than existing methods.
The model-based calculation enables efficient and comprehensive analysis of mining fairness. In this study, we analyze mining fairness using the model-based calculation. First, we examine mining fairness in a simple network composed of two miners. Subsequently, we investigate the impact of hashrate distribution on mining fairness. Finally, we reassess existing blockchain network improvement methods from the perspective of mining fairness.
2 Rounds
Figure 2: Rounds in a blockchain. Each interval enclosed by consecutive orange dotted lines represents a single round in the blockchain.
Herein, to accurately capture and incorporate the impact of forks, we introduce the concept of rounds in a blockchain. A round is a time interval defined from a global perspective. Specifically, round $r$ refers to the time from the first generation of a block at height $r$ to the first generation of a block at height $r+1$. The height of round $r$ is defined as $r$.
Because forks occur probabilistically within a blockchain network, the number of blocks per round is not always equal to one. After the first block at height $r$ has been generated, another miner may generate a new block before the first block is propagated throughout the network. It should be noted that not all blocks generated in round $r$ have a height of $r$. For example, a miner unaware of the block at height $r-1$ will generate a block at height $r-1$ during round $r$ (Fig. 2).
The round start rate of each miner is defined as the probability that the miner initiates a round. Note that the round start rate does not equal the proportion of the hashrate, owing to the occurrence of forks.
With the introduction of rounds, we can formally define the fork rate. The fork rate is the probability that the number of blocks per round is two or more. When the number of blocks per round is two or more, the blockchain diverges, or forks. Therefore, forks (and fork rates) are named as such because they function similarly to a traditional fork.
3 Model-based calculation
3.1 Model
We approximate a complex blockchain network using a simplified model to calculate mining fairness. First, we define the set of miners as $V$, and let $N$ be the number of elements in $V$. The proportion of the hashrate of miner $i$ ($i\in V$) is denoted by $\alpha_{i}$. When a new block is generated within the network, the probability that miner $i$ has generated that block is equal to the proportion of their hashrate, $\alpha_{i}$. The number of blocks generated in each round is assumed to be at most two, implying that there is at most one fork per round. We assume that the block rewards are equal.
Let $F_{ij}$ be the probability that miner $j$ generates a block that causes a fork within round $r$ started by miner $i$. After a fork occurs, as additional blocks are generated, one of the blocks will be incorporated into the main chain, while the other will not be. Let $W_{ij}$ be the probability that the block generated by miner $i$ is incorporated into the main chain under the conditions that (a) round $r$ starts with the block generated by miner $i$ and (b) miner $j$ generates a block that causes a fork.
3.2 Definition of Mining Fairness
Before calculating mining fairness, we first have to properly define it. In this study, mining fairness is divided into local mining fairness and global mining fairness. We define local mining fairness, $LF$, based on two measures as follows:
$LF_{1}(i)$ $=r_{i}-\alpha_{i},$ (1)
$LF_{2}(i)$ $=\frac{LF_{1}(i)}{\alpha_{i}},$ (2)
where $r_{i}$ refers to the block reward rate for each miner, $LF_{1}$ denotes the profit of each miner, and $LF_{2}$ denotes the profit rate of each miner.
Next, we define global mining fairness, $GF$, using local mining fairness, as follows:
$GF_{1}$ $=\sum_{i\in V}LF_{1}(i)\quad(LF_{1}(i)>0),$ (3)
$GF_{2}$ $=\max_{i\in V}LF_{2}(i)-\min_{i\in V}LF_{2}(i),$ (4)
where $GF_{1}$ is the sum of the $LF_{1}$ values that are positive, and $GF_{2}$ is the maximum difference in the profit rates. Other mining fairness measures can also be defined using $LF$.
3.3 Calculation of Mining Fairness
This section demonstrates a computational method for mining fairness based on the previously presented model. First, we determine the round start rate. Next, we calculate the local mining fairness $LF_{1}$, which is the difference between each miner’s block reward rate and the proportion of the hashrate. We also determine each miner’s profit rate $LF_{2}$. After the local mining fairness has been calculated, the global mining fairness can be easily determined.
Let $X_{r}$ be a random variable representing the miner that generates the block that starts round $r$. Then, the following equation holds:
$P(X_{r+1}=i)=\sum_{j\in V}\left(\alpha_{i}(1-F_{ji})+\sum_{k}\alpha_{k}F_{jk}\alpha_{i}\right)P(X_{r}=j).$ (5)
Note that $P(X_{r+1}=i)$ is dependent only on $P(X_{r}=j)$. Therefore, the stochastic process $\{X_{r}\}_{r=0}^{\infty}$ is a Markov chain. Additionally, because $F_{ij}$ is less than 1 and $\alpha_{i}(1-F_{ji})+\sum_{k}\alpha_{k}F_{jk}\alpha_{i}$ is usually positive, this Markov chain is ergodic in most cases. Consequently, a unique stationary distribution exists, and the limiting distribution is stationary. We can then determine the stationary distribution by iterating (5).
Let the limiting distribution be $\pi$. This represents the distribution of the miners that generate blocks that start rounds after sufficient time has passed. Using $\pi$, the block reward rate $r_{i}$ for each miner is given by the following equation:
$r_{i}=$ $\pi(i)(1-\sum_{j\in V}\alpha_{j}F_{ij}+\sum_{j\in V}\alpha_{j}F_{ij}W_{ij})$
$+\sum_{j\in V}\pi(j)\alpha_{i}F_{ji}(1-W_{ji}).$ (6)
Thus, $LF_{1}$ of miner $i$ is as follows:
$LF_{1}(i)=$ $r_{i}-\alpha_{i}$ (7)
$=$ $\pi(i)(1-\sum_{j\in V}\alpha_{j}F_{ij}+\sum_{j\in V}\alpha_{j}F_{ij}W_{ij})$
$+\sum_{j\in V}\pi(j)\alpha_{i}F_{ji}(1-W_{ji})-\alpha_{i},$ (8)
whereas $LF_{2}$ can be calculated as follows:
$LF_{2}(i)=\frac{LF_{1}(i)}{\alpha_{i}}.$ (9)
3.4 Algorithm
In this section, we describe the algorithm used in this study to calculate mining fairness, as detailed in Section 3.3.
Algorithm 1 employs an iterative method to compute the mining fairness of each miner. The round start rate calculation is performed calculated between lines 9 and 27. Specifically, The fork rate is precomputed between lines 10 and 16. The variable loop manages the operations executed in each iteration. The calculations within the for loop from lines 20 to 26 follow the same process as described in (5). Mining fairness is computed between lines 28 and 34, with the for loop calculations corresponding to equations (8) and (9).
Algorithm 1 Calculation of local mining fairness
1:The following variables are provided by the model:
2:$V$triangleright set of miners
3:$N$triangleright number of miners
4:$\alpha[N]$triangleright proportion of hashrate
5:$F[N][N]$triangleright fork rate
6:$W[N][N]$triangleright winning rate
7:
8:Our goal is to calculate the following values:
9:$\pi[N][2]$triangleright round start rate
10:$LF_{1}[N]$triangleright $LF_{1}$
11:$LF_{2}[N]$triangleright $LF_{2}$
12:
13:Calculating the round start rate of each miner:
14:$\epsilon$triangleright error
15:$dp[N]$triangleright for dynamic programming
16:for $i\in V$ do
17:     $dp[i]\leftarrow 0$
18:     for $j\in V$ do
19:         $dp[i]\leftarrow dp[i]+\alpha_{j}F[i][j]$
20:     end for
21:end for
22:$loop\leftarrow 0$
[…]
25:     for $i\in V$ do
[…]
27:         for $j\in V$ do
[…]
30:         end for
31:     end for
32:end while
33:
34:Calculating the local fairness for each miner:
35:for $i\in V$ do
[…]
37:     for $j\in V$ do
38:         $LF_{1}[i]\leftarrow LF_{1}[i]+\pi(j)\alpha_{i}F_{ji}(1-W_{ji})-\pi(i)\alpha_{j}F_{ij}(1-W_{ij})$
39:     end for
40:     $LF_{2}[i]\leftarrow LF_{1}[i]/\alpha[i]$
41:end for
3.5 Parameters
3.5.1 How to Determine $F_{ij}$
Let $T$ denote the average block generation interval, and let $T_{ij}$ represent the time it takes for a block generated by miner $i$ to be received by miner $j$. Then, $F_{ij}$ is determined as follows:
[…]
$=1-e^{-\frac{T_{ij}}{T}}.$ (11)
3.5.2 How to determine $W_{ij}$
Prior to any explanations regarding $W_{ij}$, it is first essential to discuss the concept of chain ties.
Each miner constructs chains from their blocks and selects one main chain among them. The rule for selecting this chain is known as the fork choice rule. In Bitcoin, for example, the longest chain rule, which selects the longest chain, is adopted.
However, in some cases, the fork choice rule alone may not uniquely determine the main chain owing to the occurrence of forks. This situation is known as a chain tie. To resolve a chain tie, a tie-breaking rule is implemented. In Bitcoin, for example, the first-seen rule, which selects the chain received first, is adopted.
We categorize practical tie-breaking rules as follows:
First-seen rule
Selects the earliest arriving chain among the chains in a tie. Used in Bitcoin.
Random rule
Randomly selects a chain among the chains in a tie [4]. Proposed as a countermeasure to selfish mining. Used in Ethereum.
Last-generated rule
Selects the latest chain among the chains in a tie [17, 18]. Suppresses selfish mining more effectively than the random rule.
Next, we explain how to determine $W_{ij}$. The value of $W_{ij}$ is significantly influenced by the hashrate of miners mining on the block generated by miner $i$ during a fork. More specifically, $W_{ij}$ is largely affected by the following two factors:
Tie-breaking rule
During a fork, chain ties often occur. The tie-breaking rule determines the block on which miners, other than the block generator, will mine.
Proportion of hashrate of the block generator
The block generator mines on its own generated block regardless of the tie-breaking rule.
Other factors, such as the block propagation time and the number of miners participating in the network, also influence $W_{ij}$. Section 4 provides further details on specific methods for calculating $W_{ij}$.
4 Validation
Herein, we validate the capability of our proposed model-based calculation to accurately determine mining fairness. First, we examine the assumption that the number of blocks per round is at most two from the perspective of the scale of the fork. Next, we compare the results of simulation experiments with those of the model-based calculation.
Whereas it is excessively time-consuming to calculate mining fairness via simulations of networks composed of many (approximately 100 or more) miners, it is feasible to calculate mining fairness accurately and relatively quickly for networks with a small number of miners (2–10). In this study, we perform the validation using networks consisting of two and ten miners and show that our proposed model-based calculation determines mining fairness much more accurately than existing methods.
4.1 Examining the Scale of Forks
The model-based calculation disregards the impact of large-scale forks. Specifically, it assumes that the number of blocks per round is at most two. In this study, we investigate the effects of large-scale forks.
Regarding the Scale of Forks First, we establish some facts regarding fork rates. Let the hashrate of miner $i$, where $i\in V$, be $M_{i}$. Let the total network hashrate be $M_{all}$. Additionally, let the probability of successfully generating a block with one hash calculation be $p$. Then, the average number of hash calculations required to generate a block is $1/p$. Therefore, the following equation holds:
$\frac{1}{pM_{all}}=T,$ (12)
where $T$ is the average block generation interval.
Next, let $N_{i}$ be the total number of hash calculations performed by miners who are unaware of the block generated by miner $i$. In this case, the following equation holds:
$N_{i}=\sum_{j\in V}M_{j}T_{ij}.$ (13)
Let $T_{W,i}$ be the hashrate-weighted average block propagation time for the block generated by miner $i$. Then, the following equation holds:
$T_{W,i}=\sum_{j\in V}\alpha_{j}T_{ij}.$ (14)
Therefore, from (12), (13), and (14), the following equation holds:
$pN_{i}$ $=p\sum_{j\in V}M_{j}T_{ij}$ (15)
$=\sum_{j\in V}\frac{M_{j}}{M_{all}}\frac{T_{ij}}{T}$ (16)
$=\frac{T_{W,i}}{T}.$ (17)
Next, we examine the occurrence rate of forks based on their scale. Let random variable $C_{i}$ denote the number of blocks in the round initiated by miner $i$. In this case, the following holds:
[…]
$=\sum_{j\in V}\alpha_{j}e^{-\frac{T_{ij}}{T}},$ (19)
where $P(C_{i}=1)$ denotes the probability that no forks occur. Then, the probability $P(C_{i}\neq 1)$ that a fork occurs is
$P(C_{i}\neq 1)$ $=1-P(C_{i}=1)$ (20)
$=1-\sum_{j\in V}\alpha_{j}e^{-\frac{T_{ij}}{T}}.$ (21)
The probability that the number of blocks in a round will be 3 or more satisfies the following inequality:
$P(C_{i}\geq 3)$ $\leq\sum_{k=2}^{\infty}\binom{N_{i}}{k}p^{k}(1-p)^{N_{i}-k}$ (22)
$=\sum_{k=2}^{\infty}\frac{N_{i}\cdots(N_{i}-k+1)}{k!}p^{k}(1-p)^{N_{i}-k}$ (23)
$\leq\sum_{k=2}^{\infty}\frac{(pN_{i})^{k}}{k!}e^{-p(N_{i}-k)}$ (24)
$=e^{-pN_{i}}\sum_{k=2}^{\infty}\frac{(e^{p}pN_{i})^{k}}{k!}$ (25)
$=e^{-pN_{i}}(e^{e^{p}pN_{i}}-1-e^{p}pN_{i})$ (26)
$=e^{-\frac{T_{W,i}}{T}}(e^{e^{p}\frac{T_{W,i}}{T}}-1-e^{p}\frac{T_{W,i}}{T})$ (27)
$\frac{T_{W,i}}{T}\ \text{is constant}]{p\ \to\ 0}1-(1+\frac{T_{W,i}}{T})e^{-\frac{T_{W,i}}{T}}.$ (28)
When the number of blocks in a round is 2 or more, at least 3 hash calculations succeed before all the blocks are fully shared, thus satisfying (22). Equation (27) is obtained by substituting (17) into (26).
From (28), it follows that the probability that the number of blocks in a round will be 2 satisfies the following inequality:
$P(C_{i}=2)$ $=P(C_{i}\neq 1)-P(C_{i}\geq 3)$ (29)
$\geq\sum_{j\in V}\alpha_{j}(1-e^{-\frac{T_{ij}}{T}})-\left\{1-(1+\frac{T_{W,i}}{T})e^{-\frac{T_{W,i}}{T}}\right\}$ (30)
$=(1+\frac{T_{W,i}}{T})e^{-\frac{T_{W,i}}{T}}-\sum_{j\in V}\alpha_{j}e^{-\frac{T_{ij}}{T}}.$ (31)
Impact By Fork Scale: We define the impact $I_{1}$ for rounds with one block, impact $I_{2}$ for rounds with two blocks, and impact $I_{3}$ for rounds with three or more blocks as follows:
$I_{1}$ $=e^{-\frac{d}{T}},$ (32)
$I_{2}$ $=(1+\frac{d}{T})e^{-\frac{d}{T}}-\sum_{j\in V}\alpha_{j}e^{-\frac{d}{T}}$ (33)
$=\frac{d}{T}e^{-\frac{d}{T}},$ (34)
$I_{3}$ $=1-(1+\frac{d}{T})e^{-\frac{d}{T}}.$ (35)
These definitions are obtained by substituting $T_{ij}=d$ into (19), (22), and (28). It should be noted that $I_{2}$ is defined based on the lower bound, whereas $I_{3}$ is defined based on the upper bound. In other words, $I_{2}$ is evaluated to be smaller, whereas $I_{3}$ is evaluated to be larger.
Figure 3: Comparison between $I_{3}$ and $I_{1}+I_{2}$.
Figure 4: Comparison between $I_{3}$ and $I_{2}$.
TABLE I: Influence of $I_{3}$ Relative to $I_{1}$ and $I_{2}$.
$d/T$
0.01 0.1 0.5
$I_{3}/(I_{1}+I_{2})$ 0.0000486868 0.00384871 0.0364743
$I_{3}/I_{2}$ 0.0050167084 0.0517091 0.297442
In the model-based calculation of mining fairness, we consider cases in which the number of blocks per round is two or fewer, thereby ignoring $I_{3}$. Thus, we compare $I_{3}$ with $I_{1}$ and $I_{2}$. Figs. 3 and 4 show $I_{3}/(I_{1}+I_{2})$ and $I_{3}/I_{2}$, respectively, as $d/T$ is varied from 00 to $1$. The specific numerical values are listed in Table I. It is evident that as $d/T$ decreases, the influence of $I_{3}$ diminishes. Although it is not directly demonstrated herein whether the model-based calculation can determine actual mining fairness, it can be inferred that the model-based calculation will be effective in scenarios where $d/T$ is small.
4.2 Network Composed of Two Miners
The comparison of the impact of forks by scale in Section 4.1 offers intuitive insight into how the assumption of the model affects the model-based calculation. However, it does not address how accurately the model-based calculation matches the actual numerical results for mining fairness. In this section, we validate the model-based calculation using a simple network composed of two miners in blockchain network simulations.
4.2.1 Model-Based Calculation
We perform the model-based calculation for a network comprising two miners. The calculations follow the procedure outlined in Section 3.3.
Prior to these calculations, we first provide some relevant definitions. Let the two miners in the network be $Miner_{A}$ and $Miner_{B}$. Let the proportion of the hashrate of $Miner_{A}$ be $\alpha_{A}$ and that of $Miner_{B}$ be $\alpha_{B}$, where $\alpha_{A}+\alpha_{B}=1$. Let $T$ be the average block generation interval, and $d$ be the block propagation time. Let $\pi_{A}$ and $\pi_{B}$ be the round start rates of $Miner_{A}$ and $Miner_{B}$, respectively. Next, we define $f$ as follows:
$f$ $=1-e^{-\frac{d}{T}},$ (36)
where $f$ denotes the probability that $Miner_{B}$ (or $Miner_{A}$) will create a fork when $Miner_{A}$ (or $Miner_{B}$) starts a round and the other miner generates the next block.
Next, we calculate the round start rate. Considering that a sufficiently long time has passed, the following equation holds:
$\pi_{B}=\pi_{A}(\alpha_{B}f\alpha_{B}+\alpha_{B}(1-f))+\pi_{B}(\alpha_{B}+\alpha_{A}f\alpha_{B})$ (37)
$\Rightarrow$ $\pi_{B}(1-\alpha_{B}-\alpha_{A}f\alpha_{B})=\pi_{A}(\alpha_{B}f\alpha_{B}+\alpha_{B}(1-f))$ (38)
$\Rightarrow$ $(1-\pi_{A})(1-\alpha_{B}-\alpha_{A}f\alpha_{B})=\pi_{A}(\alpha_{B}f\alpha_{B}+\alpha_{B}(1-f))$ (39)
$\Rightarrow$ $\pi_{A}(f\alpha_{B}(\alpha_{B}-1-\alpha_{A})+1)=1-\alpha_{B}-\alpha_{A}f\alpha_{B}$ (40)
$\Rightarrow$ $\pi_{A}(1-2\alpha_{A}\alpha_{B}f)=1-\alpha_{B}-\alpha_{A}f\alpha_{B}$ (41)
$\Rightarrow$ $\pi_{A}=\alpha_{A}\frac{1-\alpha_{B}f}{1-2\alpha_{A}f\alpha_{B}}.$ (42)
Because $\pi_{A}$ has already been calculated, $\pi_{B}$ can then be determined as follows:
$\pi_{B}$ $=1-\pi_{A}$ (43)
$=\alpha_{B}\frac{1-\alpha_{A}f}{1-2\alpha_{A}f\alpha_{B}}.$ (44)
Afterward, we calculate the probability $W_{AB}$ that the block generated by $Miner_{A}$ is incorporated into the main chain if $Miner_{B}$ generates a block by forking immediately after $Miner_{A}$ starts a round. For simplicity, we assume that the block of $Miner_{B}$ conflicts with the chain of $Miner_{A}$. That is, we ignore the case in which the block height of $Miner_{B}$ is smaller than that of $Miner_{A}$.
[…]
$=$ $\alpha_{A}\alpha_{A}+\alpha_{A}\alpha_{B}(1-f)$
$+2\alpha_{A}\alpha_{B}f\{\alpha_{A}\alpha_{A}+\alpha_{A}\alpha_{B}(1-f)+2\alpha_{A}\alpha_{B}f(\cdots)\}$ (46)
$=$ $\{\alpha_{A}\alpha_{A}+\alpha_{A}\alpha_{B}(1-f)\}\frac{1}{1-2\alpha_{A}\alpha_{B}f}$ (47)
$=$ $\alpha_{A}\frac{1-\alpha_{B}f}{1-2\alpha_{A}\alpha_{B}f},$ (48)
where $\alpha_{A}$ and $\alpha_{B}$ denote the probabilities of $Miner_{A}$ and $Miner_{B}$, respectively, generating a block. The superscript numbers on $\alpha_{A}$ and $\alpha_{B}$ indicate the differences in the block height from the block that initially caused the chain tie. The value of $W_{BA}$ can also be derived from $W_{AB}$ as follows:
$W_{BA}$ $=\alpha_{B}\frac{1-\alpha_{A}f}{1-2\alpha_{A}\alpha_{B}f}.$ (49)
Then, the following relationships hold:
$\pi_{A}$ $=W_{AB},$ (50)
$\pi_{B}$ $=W_{BA}.$ (51)
Using $\pi_{A}$ and $\pi_{B}$, we determine $LF_{1}(A)$ as follows:
$LF_{1}(A)=\pi_{A}+(\alpha_{A}-\alpha_{B})f\pi_{A}\pi_{B}-\alpha_{A}.$ (52)
4.2.2 Simulation Settings
Based on the blockchain network simulator SimBlock [19], we developed another event-driven simulator composed of two miners. Our simulator is capable of simulating forks of any scale, similar to those in real blockchain systems.
We examined variations in $d/T$, i.e., the ratio of the block propagation time to the average block generation interval, with values of $0.1$, $0.3$, and $0.5$. The block propagation time was kept constant among all miners.
We also examined variations in $\alpha_{A}$, i.e., the proportion of hashrate of miner $A$, with values of $0.1$ and $0.3$. A value of $0.5$ was not considered because, in this case, mining fairness is completely maintained owing to the symmetry of the network.
Each simulation consisted of 10 billion rounds.
TABLE II: Errors Between Simulation and Model-Based Calculation Results for $LF_{1}$.
$d/T$
0.1 0.3 0.5
$\alpha_{A}$ 0.3 0.00059936 0.00792584 0.0203151
0.1 0.000315594 0.0031603 0.00744887
4.2.3 Validation Results
The errors between the simulation and model-based calculation results are listed in Table II. Here, error is defined as the relative error as follows:
$\frac{d_{euclid}(LF_{simulation},LF_{MBC})}{d_{euclid}(LF_{simulation,0})},$ (53)
where $d$ is the Euclidean distance, $LF_{simulation}$ is the vector of the simulated values of local mining fairness for each miner, and $LF_{MBC}$ is the vector of the model-based calculated values of local mining fairness.
Based on the error values, it can be seen that the model-based calculation can compute mining fairness with high accuracy. It can also be observed that the accuracy deteriorates as $d/T$ increases. This is likely because, as seen in Section 4.1, the impact of having more than three blocks per round becomes more significant as $d/T$ increases.
4.3 Network Composed of 10 Miners
In this section, we validate the model-based calculation of mining fairness on a network composed of 10 miners. Compared to a network with two miners, a network with ten miners introduces additional elements such as tie-breaking rules, hashrate distribution, and block propagation time. This allows us to demonstrate that the proposed model-based calculation method is effective even in more complex networks.
4.3.1 $W_{ij}$ in a Network Composed of Multiple Miners
In a network with more than two miners, it is necessary to consider tie-breaking rules. Here, we demonstrate how to determine $W_{ij}$ for a network with multiple miners according to different tie-breaking rules. It is assumed that all forks cause chain ties.
First-Seen Rule: We assume that miner $i$ starts a round, and then miner $j$ causes a chain tie in the same round. Let $p_{i,j,k}$ be the probability that miner $k$ mines on the block generated by miner $i$. The time $T_{ij}$ it takes for the block generated by miner $i$ to reach miner $j$ is assumed to be a fixed value that depends only on $i$ and $j$.
When $T_{ik}<T_{jk}$, regardless of the time when miner $j$ generates the block, the block generated by miner $i$ will reach miner $k$ first, and hence, $p_{i,j,k}=1$. Similarly, when $T_{ik}<T_{ij}+T_{jk}$, the block generated by miner $j$ will reach miner $k$ first, and hence, $p_{i,j,k}=0$. In other cases, the following equation holds:
[…]
$=\frac{e^{-\frac{T_{ik}-T_{jk}}{T}}-e^{-\frac{T_{ij}}{T}}}{1-e^{-\frac{T_{ij}}{T}}}.$ (55)
Equation (54) defines the probability that the block of miner $i$ reaches miner $k$ first under the condition that a chain tie occurs. Equation (55) substitutes $F_{ij}$ into (54) based on (11).
From $p_{i,j,k}$, the value of $W_{ij}$ is determined as follows:
$W_{ij}$ $=\sum_{k\in V}\alpha_{k}p_{i,j,k}.$ (56)
Random Rule: Herein, mining is performed by selecting a block randomly during a chain tie. The value of $W_{ij}$ is given by the following equation:
$W_{ij}$ $=\alpha_{i}+\frac{1-\alpha_{i}-\alpha_{j}}{2}.$ (57)
Last-Generated Rule: In this rule, mining is performed by selecting the most recently generated block during a chain tie. The value of $W_{ij}$ is given by the following equation:
$W_{ij}$ $=\alpha_{i}.$ (58)
Figure 5: Hashrate distribution settings used in validation of model-based calculation.
4.3.2 Simulation Settings
The simulator used in this validation was an extended version of a network simulator composed of two miners. The number of miners was set to 10. The hashrate distribution was based on that of Bitcoin [20]. The hashrate distribution settings are shown in Fig. 5.
In this validation, the ratios of the average block propagation time to the average block generation interval, $d/T$, were varied among the values of $0.01$, $0.04$, $0.07$, and $0.1$. These settings cover most blockchain systems; for instance, in the case of Bitcoin, $d/T$ is approximately equal to $0.00576$, whereas in the case of Ethereum, $d/T=0.068$ [5, 21]. The block propagation time distribution among the different miners followed an exponential distribution [2], whereas the block propagation time to oneself was set to 0. Additionally, the previously described tie-breaking rules, i.e., the first-seen rule, random rule, and last-generated rule, were examined.
Each simulation consisted of 10 billion rounds.
TABLE III: Mean and Standard Deviation (SD) of Errors Between Simulation and Model-Based Calculation Results for $d/T=0.01$ and $0.04$ for Network of Ten Miners.
First-seen rule Random rule Last-generated rule
d/T
0.01 0.04 0.01 0.04 0.01 0.04
Round start rate Mean 0.0000201254 0.000132747 0.0000208543 0.000134614 0.0000207662 0.000135
SD 0.00000639285 0.0000648318 0.00000663684 0.0000664632 0.00000700959 0.0000649111
$LF_{1}$ Mean 0.00891706 0.0221225 0.0153555 0.0507862 0.0214307 0.080677
SD 0.00356624 0.0115312 0.00771275 0.0289467 0.0130309 0.0613974
$LF_{2}$ Mean 0.0108707 0.0209106 0.0164265 0.0439469 0.0210495 0.0702858
SD 0.00373994 0.00828696 0.00546804 0.0146293 0.0110262 0.0426435
TABLE IV: Mean and Standard Deviation (SD) of Errors Between Simulation and Model-Based Calculation Results for $d/T=0.07$ and $0.1$ for Network of Ten Miners.
First-seen rule Random rule Last-generated rule
d/T
0.07 0.1 0.07 0.1 0.07 0.1
Round start rate Mean 0.000406944 0.0008385 0.0004072 0.000840025 0.000408239 0.000839217
SD 0.000215053 0.000462696 0.000215627 0.000461556 0.000217068 0.000460759
$LF_{1}$ Mean 0.0851189 0.0553571 0.0851189 0.117411 0.13679 0.188992
SD 0.0467998 0.0293375 0.0467998 0.0630781 0.104206 0.141844
$LF_{2}$ Mean 0.0737303 0.0520755, 0.0737303 0.10182 0.117051 0.159697
SD 0.0242187 0.0205341 0.0242187 0.0328015 0.0670573 0.0871817
TABLE V: Proposed Model-Based Calculation of $LF_{1}$ for Network of Ten Miners vs. Existing Method.
$d/T$
0.01 0.04 0.07 0.1
Fist-seen rule Proposed method 0.00891706 0.0221225 0.0386387 0.0553571
Exiting method 1.23734 1.23344 1.22872 1.22402
Random rule Proposed method 0.0153555 0.0507862 0.0851189 0.117411
Exiting method 1.69555 1.66842 1.64136 1.61529
Last-generated rule Proposed method 0.0214307 0.080677 0.13679 0.188992
Exiting method 1.56918 1.59359 1.60905 1.6178
TABLE VI: Proposed Model-Based Calculation of $LF_{2}$ for Network of Ten Miners vs. Existing Method.
$d/T$
0.01 0.04 0.07 0.1
Fist-seen rule Proposed method 0.0108707 0.0209106 0.0365199 0.0520755
Exiting method 0.913274 0.910973 0.908631 0.906371
Random rule Proposed method 0.0164265 0.0439469 0.0737303 0.10182
Exiting method 1.11546 1.10628 1.09711 1.08832
Last-generated rule Proposed method 0.0210495 0.0702858 0.117051 0.159697
Exiting method 1.1644 1.16133 1.15597 1.15099
4.3.3 Validation Results
We conducted 50 simulation experiments for each validation target. The errors between the simulation and model-based calculation results are shown in Tables III and IV. Here, error is defined as the relative error as in (53). The difference from the previous validation is that the number of elements in the $LF$ vector is changed from two to ten, and we investigate not only the $LF$ vector but also the round-start-rate vector.
First, we examine the round start rate. It can be seen that the model-based calculations match the simulation results with high accuracy under all conditions.
Next, we examine mining fairness. From the tables, it can be seen that the model-based calculation method is able to compute mining fairness with high accuracy. It can also be observed that the accuracy deteriorates as $d/T$ increases.
Furthermore, it is found that the accuracy of the mining fairness calculations is not as high as that of the round-start-rate calculations or of the mining fairness calculations for a network composed of two miners. This is because the mining fairness calculation requires $W_{ij}$.
Additionally, it is observed that the first-seen rule is more accurate than the random rule or the last-generated rule. This finding indicates that the calculation of $W_{ij}$ based on the first-seen rule is superior. This is because the influence of blocks up to the second one in a given round is stronger under the first-seen rule.
We also compared our proposed model-based calculation with a state-of-the-art method. This comparison method is the same as that proposed except that the latter considers the round start rate [13]. By contrast, the comparison method automatically assumes that the round start rate is equal to the hashrate. Tables V and VI present the results. As can be seen, our model-based calculation significantly improves the accuracy compared with that of the state-of-the-art method. This result demonstrates the importance of considering the impact of forks on the round start rate.
5 Analysis
In this section, we present a detailed analysis of mining fairness using the model-based calculation. First, we analyze mining fairness for a simple network composed of two miners (Section 5.1). Next, we examine the impact of hashrate distribution on mining fairness (Section 5.2). Finally, we reassess existing blockchain network improvement methods from the perspective of mining fairness (Section 5.3).
5.1 Network Composed of Two Miners
In this section, we provide a thorough examination of mining fairness in a simple network that includes only two miners. Analyzing such a network offers two advantages: there is no need to consider tie-breaking rules, and it allows for a more general analysis without substituting specific values for the parameters. We investigate the properties of mining fairness in this network using the proposed model-based calculation.
5.1.1 Analysis Subject
We analyze a network composed of two miners ($Miner_{A}$ and $Miner_{B}$). The block propagation time $d$ is fixed to a constant value among all miners. The block propagation time to itself is assumed to be 0.
5.1.2 Analysis Results
We conducted an analysis based on the model-based calculation using $W_{ij}$ (see Section 4.2), which was validated in a network composed of two miners.
Figure 6: $LF_{1}(A)$ for network composed of two miners.
TABLE VII: $LF_{1}$ for Different Values of $d/T$ and $\alpha_{A}$.
$d/T$
0.1 0.2 0.3 0.4 0.5
$\alpha_{A}$ 0.5 0 0 0 0 0
0.7 0.0161907 0.0312103 0.045127 0.0580053 0.0699067
0.9 0.0133946 0.0249514 0.034909 0.0434791 0.0508486
First, we examine the local mining fairness $LF_{1}(A)$ for $Miner_{A}$. Table VII and Fig. 6 illustrate $LF_{1}(A)$ with variations in $d/T$, i.e., the ratio of the block propagation time to the average block generation interval, and in $\alpha_{A}$, i.e., the proportion of hashrate of $Miner_{A}$. It is evident herein that as $d/T$ increases, mining fairness deteriorates.
TABLE VIII: Maximum $LF_{1}$ and Corresponding $\alpha_{A}$ for Various Values of $d/T$.
$d/T$
0.1 0.2 0.3 0.4 0.5
$LF_{1}$ 0.0183165 0.0349033 0.0499354 0.0635683 0.0759402
$\alpha_{A}$ 0.783959 0.779441 0.775129 0.771029 0.767144
Next, we focus on $\alpha_{A}$. Considering symmetry, we perform analyses with $\alpha_{A}$ varying from 0.5 to 1, by which $LF_{1}$ is observed to concave upward. This is because as $\alpha_{A}$ increases, $Miner_{A}$ gains power to take rewards from $Miner_{B}$, but the number of blocks generated by $Miner_{B}$ decreases accordingly.
Then, we investigate the maximum value of $LF_{1}$ and the corresponding $\alpha_{A}$ for various values of $d/T$ using a ternary search. Table VIII shows the results. It can be observed that as $d/T$ increases, the maximum value of $LF_{1}$ increases, whereas $\alpha_{A}$ at which this maximum value is achieved decreases.
Next, we examine the local mining fairness of $Miner_{A}$, i.e., $LF_{2}(A)$. Table IX and Fig. 7 show $LF_{2}(A)$ with respect to variations in $d/T$ and $\alpha_{A}$. It is evident that as $d/T$ increases, mining fairness deteriorates.
Afterward, we focus on $\alpha_{A}$. Unlike $LF_{1}$, $LF_{2}$ does not exhibit any symmetry. When $\alpha_{A}$ is varied from 0.0 to 0.5, $LF_{2}$ increases monotonically. When $\alpha_{A}$ is varied from 0.5 to 1, $LF_{2}$ concaves upwards. Therefore, we investigate the maximum value of $LF_{2}$ and the corresponding $\alpha_{A}$ for various values of $d/T$ using a ternary search. Table X shows the results. It can be observed that as $d/T$ increases, the maximum value of $LF_{2}$ increases, whereas $\alpha_{A}$ at which this maximum value is achieved decreases. It is also observed that $\alpha_{A}$ at which the maximum value of $LF_{2}$ is achieved is slightly smaller than that for $LF_{1}$.
Figure 7: $LF_{2}(A)$ for network composed of two miners.
TABLE IX: $LF_{2}$ for Different Values of $d/T$ and $\alpha_{A}$.
$d/T$
0.1 0.2 0.3 0.4 0.5
$\alpha_{A}$ 0.1 -0.00133946 -0.00249514 -0.0034909 -0.00434791 -0.00508486
0.3 -0.00053969 -0.00104034 -0.00150423 -0.00193351 -0.00233022
0.5 0 0 0 0 0
0.7 0.000231296 0.000445862 0.000644671 0.000828646 0.000998668
0.9 0.000148828 0.000277238 0.000277238 0.000483101 0.000564984
TABLE X: Maximum $LF_{2}$ and Corresponding $\alpha_{A}$ for Various Values of $d/T$.
$d/T$
0.1 0.2 0.3 0.4 0.5
$LF_{1}$ 0.0239401 0.0458883 0.0660208 0.0844949 0.101452
$\alpha_{A}$ 0.74543 0.741098 0.737005 0.73315 0.729529
Finally, we examine global mining fairness. For a network composed of two miners, the property $GF_{1}$ can be understood based on $LF_{1}$. Therefore, we focus only on $GF_{2}$. Table XI and Fig. 8 show $GF_{2}$ with respect to variations in $d/T$ and $\alpha_{A}$. It can be observed that as $d/T$ increases, mining fairness deteriorates. Then, we focus on $\alpha_{A}$. Owing to symmetry, we consider the case in which $\alpha_{A}$ is between 0.5 and 1. It is observed that $GF_{2}$ increases monotonically as $\alpha_{A}$ increases.
Figure 8: $GF_{2}(A)$ for network composed of two miners.
TABLE XI: $GF_{2}$ for Different Values of $d/T$ and $\alpha_{A}$.
$d/T$
0.1 0.2 0.3 0.4 0.5
$\alpha_{A}$ 0.5 0 0 0 0 0
0.7 0.0542003 0.10448 0.151068 0.194179 0.234021
0.9 0.134094 0.249791 0.349478 0.435274 0.509051
5.2 Impact of Hashrate Distribution
In this section, we explore how the hashrate distribution affects mining fairness. Different hashrate distributions can result in varying degrees of mining fairness. For example, it is expected that the revenues of miners with larger hashrates will tend to be higher. This is because the blocks of these miners are more likely to be included in the main chain in the event of ties. We aim to analyze such patterns thoroughly by examining various hashrate distribution scenarios and their impacts on revenue distribution.
5.2.1 Analysis Subjects
Block Propagation Time and Average Block Generation Interval: We perform our analysis for various ratios of the propagation time to the average block generation interval, $d/T$, namely, 0.01, 0.04, and 0.07. The block propagation time is fixed to a constant value among all miners, and the propagation time to oneself is set to 0.
Tie-Breaking Rule: We analyze three tie-breaking rules: the first-seen rule, the random rule, and the last-generated rule.
(a) Bitcoin
(b) Dogecoin
(c) Bitcoin Cash
(d) Litecoin
(e) Ethereurm Classic
(f) Bitcoin SV
Figure 9: Hashrate distributions.
Hashrate Distribution and Number of Miners The number of miners is set to 100. We analyze six distributions of the hashrate ratio as shown in Fig. 9 [20]. Specifically, we first allocate the hashrate ratios of the Pools, excluding Others, to the miners. Then, the remaining hashrate ratio is evenly distributed among the remaining miners, ensuring that the total hashrate ratio of these miners equals the hashrate ratio of Others. The value for Others corresponds to the total hashrate ratio of solo miners.
Figure 10: $LF_{2}$ with hashrate distribution of Bitcoin under first-seen rule.
5.2.2 Analysis Results
We conducted an analysis based on the model-based calculation using $W_{ij}$ for a network composed of ten miners (see Section 4.3).
First, we confirmed that $LF_{2}$ becomes a linear function of $\alpha$. Fig. 10 shows scatter plots of $LF_{2}$ with the hashrate distribution of Bitcoin under the first-seen rule. It is evident that $(\alpha,LF_{2})$ lies on a straight line regardless of the conditions. This is also true for any analysis subject, based on the correlation coefficient, which averages 0.999995, with a standard deviation of 0.00000667953. Then, using the least-squares method, we determined the slope of each line. Table XII shows the results, and Table XIII presents the mean, standard deviation, and coefficient of variation of the slopes for each column. From the latter, it can be observed that the hashrate distribution has little effect on the slopes.
Afterward, we confirm the correlation between the slopes of $LF_{2}$ and $d/T$. First, from (21) and by fixing $d/T$, we derive the following approximate formula for the fork rate:
$P(C_{i}\neq 1)$ $=\sum_{j\in V}\alpha_{j}(1-e^{-\frac{T_{ij}}{T}})$ (59)
$=\sum_{j\in V}\alpha_{j}(1-e^{-\frac{d}{T}})$ (60)
$=(1-e^{-\frac{d}{T}}).$ (61)
Next, we compare the ratio of the fork rates and the ratio of the slopes of $LF_{2}$ when $d/T$ is $0.01$ and $0.04$, and when $d/T$ is $0.01$ and $0.07$. Table XIV presents the results. It is clear that the fork rate and slope of $LF_{2}$ have a proportional relationship. As the fork rate increases with increasing $d/T$, it can be seen that the slope of $LF_{2}$ also increases with increasing $d/T$. This indicates that mining fairness is increasingly compromised as $d/T$ increases.
TABLE XII: Comparison of Slopes of $LF_{2}$.
First-seen rule Random rule Last-generated rule $\Sigma_{i\in V}\alpha_{i}^{2}$
$d/T$
0.01 0.04 0.07 0.01 0.04 0.07 0.01 0.04 0.07
Bitcoin 0.0198907 0.0782723 0.134769 0.0198534 0.0776913 0.133039 0.019816 0.0771103 0.131308 0.171452
Dogecoin 0.0198916 0.0782865 0.134811 0.0198523 0.0776752 0.132992 0.0198131 0.0770643 0.131172 0.147123
Bitcoin Cash 0.0199034 0.0784709 0.135362 0.0198718 0.0779764 0.133884 0.0198401 0.0774819 0.132406 0.164704
Litecoin 0.0198894 0.0782515 0.134708 0.0198506 0.0776484 0.132912 0.0198118 0.0770452 0.131116 0.163495
Ethereum Classic 0.019889 0.0782462 0.134692 0.0198622 0.0778276 0.133442 0.0198353 0.077409 0.132192 0.28605
Bitcoin SV 0.0198645 0.0778646 0.133555 0.0198388 0.0774643 0.132364 0.019813 0.0770641 0.131172 0.420785
TABLE XIII: Mean, Standard Deviation (SD), and Coefficient of Variation (CV) of Slopes.
First-seen rule Random rule Last-generated rule
$d/T$
0.01 0.04 0.07 0.01 0.04 0.07 0.01 0.04 0.07
Mean 0.0198881 0.078232 0.134649 0.0198548 0.0777138 0.133105 0.0198215 0.0771958 0.131561
SD 0.000011638 0.000181306 0.000540243 0.00001021106 0.000158341 0.000469727 0.0000115562 0.000178855 0.000528655
CV 0.000585178 0.00231754 0.00401221 0.000514285 0.00203749 0.00352898 0.000583015 0.0023169 0.00401833
TABLE XIV: Comparison of Fork Rate and $LF_{2}$.
Fork rate First-seen rule Random rule Last-generated rule
($d/T$ = 0.04) / ($d/T$ = 0.01) 3.94069 3.9336 3.91409 3.89453
($d/T$ = 0.07) / ($d/T$ = 0.01) 6.79447 6.77035 6.70392 6.63727
Next, we examine the root $\alpha_{0}$ of $LF_{2}$. First, given that $LF_{2}$ can be expressed as a linear function of the proportion of hashrate $\alpha$, the following equation holds:
$LF_{2}(i)$ $=k(\alpha_{i}-\alpha_{0}),$ (62)
where $k$ is a constant determined by the tie-breaking rule and $d/T$.
From (62), $LF_{1}$ is derived as follows:
$LF_{2}(i)\alpha_{i}$ $=\alpha_{i}k(\alpha_{i}-\alpha_{0})$ (63)
$\iff LF_{1}(i)$ $=k\alpha_{i}^{2}-k\alpha_{0}\alpha_{i}.$ (64)
By summing $LF_{1}$, we obtain
$\sum_{i\in V}LF_{1}(i)$ $=\sum_{i\in V}\left(k\alpha_{i}^{2}-k\alpha_{0}\alpha_{i}\right)$ (65)
$\iff 0$ $=k\sum_{i\in V}\alpha_{i}^{2}-k\alpha_{0}\sum_{i\in V}\alpha_{i}$ (66)
$\iff\alpha_{0}$ $=\sum_{i\in V}\alpha_{i}^{2}.$ (67)
Therefore, $\alpha_{0}$ is equal to $\sum_{i\in V}\alpha_{i}^{2}$. It is evident from the following equation that $\sum_{i\in V}\alpha_{i}^{2}$ is related to the variance in the proportion of hashrate:
$\sum_{i\in V}\alpha_{i}^{2}$ $=N\left(\text{Var}[\alpha]+\frac{1}{N^{2}}\right).$ (68)
Therefore, given that $N$ is fixed at 100, as the variance increases, $\sum_{i\in V}\alpha_{i}^{2}$ also increases, which in turn increases the root of $LF_{2}$. This implies that it becomes more difficult to achieve a high-profit rate.
TABLE XV: $GF_{1}$ and $GF_{2}$ for $d/T=0.01$.
first-seen rule random rule last-generated rule Variance
$GF_{1}$ $GF_{2}$ $GF_{1}$ $GF_{2}$ $GF_{1}$ $GF_{2}$
Bitcoin 0.000833972 0.00539058 0.000832568 0.0053808 0.000831165 0.00537101 0.00161452
Dogecoin 0.000694512 0.00489436 0.000693225 0.00488519 0.000691937 0.00487602 0.00137123
Bitcoin Cash 0.00145547 0.00725821 0.00145333 0.00724673 0.0014512 0.00723525 0.00154704
Litecoin 0.000591993 0.00491319 0.000590934 0.00490398 0.000589874 0.00489477 0.00153495
Ethereum Classic 0.00197496 0.00972674 0.00197285 0.00971435 0.00197075 0.00970196 0.0027605
Bitcoin SV 0.000719277 0.00981355 0.000718751 0.00980117 0.000718225 0.00978879 0.00410785
Figure 11: $GF_{1}$ for various hashrate distribtutions.
Figure 12: $GF_{2}$ for various hashrate distribtutions.
Next, we examine the relationship between the variance and global mining fairness in detail. We consider only the case in which $d/T=0.01$ under the first-seen rule. This is because, from the analysis of the slope of $LF_{2}$ and $d/T$ mentioned earlier, if the result for $d/T=0.01$ is understood, the results for $d/T=0.04$ and $0.07$ can also be inferred. Additionally, Table XII shows that the impact of tie-breaking rules on mining fairness is significantly smaller than the effect of $d/T$. Table XV and Figs. 11 and 12 show the global mining fairness and how it changes with respect to the hashrate distribution. It can be observed that $GF_{2}$ increases as the variance increases. A similar trend is observed for $GF_{1}$. However, in the case of Bitcoin SV, where the variance is large, $GF_{1}$ has a small value. This is because the root of $LF_{2}$ increases, resulting in a decrease in the profitability of miners with a large share of the hashrate.
5.3 Reassessing Existing Blockchain Network Improvement Methods
Despite the significance of analyzing mining fairness, existing methods for analyzing and improving blockchain networks have not been evaluated from the perspective of mining fairness. Other metrics such as fork rate, stale rate, X percentile block propagation time, communication overhead, and average propagation time cannot be used as substitutes for mining fairness. For instance, even if the overall stale rate is high, mining fairness can be achieved if all the miners have equal stale rates.
In this section, we reevaluate existing methods for improving blockchain networks from the perspective of mining fairness. We also examine the relationship between block propagation time and mining fairness. Furthermore, we analyze the partial application of these methods to improve blockchain networks.
5.3.1 Analysis Subjects
The analysis subjects are as follows:
Normal Block Relay 2019 (Normal 2019)
A block relay protocol that was widely used in Bitcoin. Herein, a node that has received a block sends an inv message to neighboring nodes to inform them of the new block. Nodes that are unaware of the block request for it by sending getdata messages. Upon receiving the getdata message, the node sends the block. The network parameter settings (network latency and bandwidth) are the same as those in 2019 [22].
Normal Block Relay 2015 (Normal 2015)
Almost the same as Normal 2019 except for the network configuration. Network parameter settings (network latency and bandwidth) are the same as those in 2015 [22]. The purpose of including this network improvement method in the analysis is to investigate the effects of network development on mining fairness.
Compact Block Relay 2019 (Compact 2019)
A block relay protocol that is intended to reduce traffic [23]. Its concept is very simple: The nodes send compact blocks instead of full blocks. Compact blocks contain short transaction IDs, not raw transactions. When nodes receive compact blocks, they attempt to reconstruct the full blocks. If the reconstruction fails, the nodes require the transactions corresponding to the short transaction IDs. Compact Block Relay has two modes: high-bandwidth mode and low-bandwidth mode. Nodes using the high-bandwidth mode send compact blocks without sending inv messages and verifying the blocks. Here, we consider only the low-bandwidth mode. The network parameter settings (network latency and bandwidth) are the same as those in 2019.
Graphene 2019
A block relay protocol that enables lower traffic consumption than that for Compact Block Relay [24]. Graphene uses Bloom filters [25] and invertible Bloom lookup tables (IBLTs) [26] instead of full blocks. The success rate of decoding IBLTs is very high when proper parameters are set, and the sizes of IBLTs are lower than those of compact blocks. The specific protocol and implementation of Graphene are very complex. Fortunately, a study by Muhammad et al. provides an entire protocol that is easy to understand [27]. The network parameter settings (network latency and bandwidth) are the same as those in 2019.
Relay Network 2019 (Relay 2019)
A network for fast block propagation. Nodes can inform blocks throughout the blockchain network by using a relay network. There are several types of relay networks. The type of relay networks we consider is the same as those investigated by Otsuki et al [28]. This type of relay network conforms to the basic structure of relay networks. When nodes generate blocks, they utilize a relay server to send those blocks. The relay servers then share the blocks with other relay servers and clients. The network parameter settings (network latency and bandwidth) are the same as those in 2019.
The calculation of mining fairness also accounts for tie-breaking rules. As in the previous analysis, we perform our analyses under the first-seen rule, random rule, and last-generated rule.
5.3.2 Analysis Method and Setup
Figure 13: Combining blockchain network simulations and model-based calculation.
The analysis was conducted by combining blockchain network simulations and the model-based calculation. First, we simulated the blockchain network improvement methods. Subsequently, from these simulations, we obtained the proportion of hashrate for each miner and the block propagation time. Using the simulation results, we performed the model-based calculation to determine mining fairness (Fig. 13).
We used SimBlock, a blockchain network simulator [19], to determine the block propagation time for each method. In SimBlock, miners had attributes such as region, hashrate, and neighbor nodes. There were a total of six regions. The bandwidth and latency were determined between the regions. The hashrate distribution followed a normal distribution with a mean of $4\cdot 10^{8}$ (/s) and a standard deviation of $1\cdot 10^{8}$ (/s). Each miner had eight outbound connections and up to 125 inbound connections. The number of miners was set to 1000, and the average block generation interval, $T$, was set to 600 s.
For Normal 2019, Normal 2015, Compact 2019, and Graphene 2019, the implementations were the same as those provided by Muhammad et al [27]. For Relay 2019, the implementation was referenced from that of Otsuki et al [28]. Specifically, relay servers were newly installed in each region. The bandwidth for the communication between the relay servers was reduced to 1/10. Additionally, the relay servers propagated blocks in parallel with the relay clients within the same region. The relay clients sent blocks to the relay servers only when the clients generated them.
We also set the parameters for each block size (full blocks, compact blocks, and graphene blocks) and the success rate of reconstructing full blocks in the case of Compact Block Relay and Graphene using log data provided by Muhammad et al [27, 29].
For the block propagation time between miners, each miner generated and propagated blocks 10 times, and the average propagation time was calculated.
5.3.3 Analysis Results
We analyze the aforementioned methods from the perspective of mining fairness. First, we investigate global mining fairness and the relationships between mining fairness, block propagation time, and fork rate. Afterward, we analyze the effect of partial application on mining fairness.
Before analyzing global mining fairness, we first define the hashrate-weighted average block propagation time, $T_{W}$, which we will use in the subsequent analyses:
$T_{W}=\sum_{i\in V}\pi(i)T_{W,i}.$ (69)
From (21), the following equation holds:
$\sum_{i\in V}\pi(i)P(C_{i}\neq 1)$ $=\sum_{i\in V}\pi(i)\sum_{j\in V}\alpha_{j}(1-e^{-\frac{T_{ij}}{T}})$ (70)
$\approx\sum_{i\in V}\pi(i)\sum_{j\in V}\alpha_{j}\frac{T_{ij}}{T}$ (71)
$=\sum_{i\in V}\pi(i)\frac{T_{W,i}}{T}$ (72)
$=\frac{T_{W}}{T}.$ (73)
That is, if $T_{ij}/T$ is sufficiently small, then $T_{W}$ is proportional to the fork rate. Therefore, by comparing $T_{W}$ and mining fairness, we can simultaneously analyze the block propagation time, fork rate, and mining fairness.
TABLE XVI: Mining Fairness for Various Methods.
First-seen rule Random rule Last-generated rule $T_{W}$(ms)
$GF_{1}$ $GF_{2}$ $GF_{1}$ $GF_{2}$ $GF_{1}$ $GF_{2}$
Normal 2019 0.00494648 0.0537581 0.0023965 0.0261654 0.00279892 0.0355703 2969.62
Normal 2015 0.0185705 0.152324 0.00859747 0.0695247 0.0127475 0.0847132 8073.2
Compact 2019 0.00243168 0.0233093 0.001209 0.0115734 0.00136458 0.0847132 685.085
Graphene 2019 0.00242575 0.0233325 0.00121145 0.0116221 0.0013509 0.0116336 604.749
Relay 2019 0.00497872 0.0771632 0.0026799 0.0425951 0.00392692 0.0575781 1377.98
Figure 14: $GF_{1}$ and $T_{W}$ for various methods.
Figure 15: $GF_{2}$ and $T_{W}$ for various methods.
The experimental results are shown in Figs. 14 and 15, which illustrate the relationship between $T_{W}$ and mining fairness, and Table XVI, which details the mining fairness for the various methods. The correlation coefficients between $T_{W}$ and mining fairness are listed in Table XVII. From Figs. 14 and 15 and Table XVII, it is evident that there is a strong correlation between $T_{W}$ and mining fairness. However, it is also clear that Normal 2019 yields better mining fairness than that offered by Relay 2019, despite the larger $T_{W}$ of the former. This implies that a smaller $T_{W}$ does not necessarily lead to an improvement in mining fairness. This highlights the significance of using mining fairness as an evaluation criterion in addition to block propagation time and fork rate.
TABLE XVII: Correlation Coefficients Between $T_{W}$ and Mining Fairness.
First-seen rule Random rule Last-generated rule
$GF_{1}$ 0.982311 0.978016 0.966078
$GF_{2}$ 0.934586 0.887302 0.852321
Next, we analyze mining fairness in partial application. We investigate the differences in benefits between miners using these methods and those who do not. The analysis subjects are Compact 2019, Graphene 2019, and Relay 2019. Before the analysis, we define partial mining fairness ($PF$) as follows:
$PF_{1}(W)$ $=\sum_{i\in W}LF_{1}(i)\quad(W\subseteq V),$ (74)
$PF_{2}(W)$ $=\max_{i\in W}LF_{2}(i)-\min_{i\in W}LF_{2}(i).$ (75)
Partial mining fairness $PF_{1}$ indicates how much benefit a group of miners $W$ gains relative to the whole, whereas $PF_{2}$ indicates the disparity in profit rates within the group of miners.
TABLE XVIII: Mining Fairness in Partial Application.
First-seen rule Random rule Last-generated rule
$PF_{1}$ $PF_{2}$ $PF_{1}$ $PF_{2}$ $PF_{1}$ $PF_{2}$
Compact 2019 $V^{+}$ 0.001 750 40.00175040.001\,750\,40.001 750 4 0.057 463 30.05746330.057\,463\,30.057 463 3 0.000 873 4380.0008734380.000\,873\,4380.000 873 438 0.029 012 80.02901280.029\,012\,80.029 012 8 0.001 205 250.001205250.001\,205\,250.001 205 25 0.041 359 70.04135970.041\,359\,70.041 359 7
$V^{-}$ 0.001 750 4-0.0017504-0.001\,750\,4- 0.001 750 4 0.058 563 60.05856360.058\,563\,60.058 563 6 0.000 873 438-0.000873438-0.000\,873\,438- 0.000 873 438 0.029 132 60.02913260.029\,132\,60.029 132 6 0.001 205 25-0.00120525-0.001\,205\,25- 0.001 205 25 0.046 181 50.04618150.046\,181\,50.046 181 5
Graphene 2019 $V^{+}$ 0.001 802 780.001802780.001\,802\,780.001 802 78 0.057 582 20.05758220.057\,582\,20.057 582 2 0.000 901 7170.0009017170.000\,901\,7170.000 901 717 0.029 134 60.02913460.029\,134\,60.029 134 6 0.001 242 810.001242810.001\,242\,810.001 242 81 0.041 380 70.04138070.041\,380\,70.041 380 7
$V^{-}$ 0.001 802 78-0.00180278-0.001\,802\,78- 0.001 802 78 0.059 779 80.05977980.059\,779\,80.059 779 8 0.000 901 717-0.000901717-0.000\,901\,717- 0.000 901 717 0.029 8050.0298050.029\,8050.029 805 0.001 242 81-0.00124281-0.001\,242\,81- 0.001 242 81 0.047 549 70.04754970.047\,549\,70.047 549 7
Relay 2019 $V^{+}$ 0.005 693 860.005693860.005\,693\,860.005 693 86 0.069 809 40.06980940.069\,809\,40.069 809 4 0.002 884 490.002884490.002\,884\,490.002 884 49 0.036 483 90.03648390.036\,483\,90.036 483 9 0.005 245 540.005245540.005\,245\,540.005 245 54 0.049 7190.0497190.049\,7190.049 719
$V^{-}$ 0.005 693 86-0.00569386-0.005\,693\,86- 0.005 693 86 0.065 393 20.06539320.065\,393\,20.065 393 2 0.002 884 49-0.00288449-0.002\,884\,49- 0.002 884 49 0.032 742 70.03274270.032\,742\,70.032 742 7 0.005 245 54-0.00524554-0.005\,245\,54- 0.005 245 54 0.051 004 80.05100480.051\,004\,80.051 004 8
Figure 16: $PF_{1}$ for various methods.
We divided the miners into two equal groups: one using a given method and the other not. Let $V^{+}$ be the set of miners using this method, and $V^{-}$ be the set of miners not using the method. Table XVIII and Fig. 16 show mining fairness in partial application. First, we focus on $PF_{1}$. It can be observed that miners who use a relay network benefit more than those who use Compact Block Relay and Graphene. This is believed to be due to the relay server propagating blocks faster only when the relay client generates a block. Additionally, based on $PF_{2}$, it is evident that for all methods, there is no significant disparity between miners using a given method and those that do not use it.
6 Discussion
6.1 Related Work
Mining Fairness for Unintentional Forks: Croman et al. proposed several metrics and solutions to address the scalability problems of blockchains [9]. They highlighted mining fairness as a potentially more valuable metric, despite its difficulty in measurement.
Since then, numerous studies have attempted to analyze mining fairness [10, 11, 12, 13, 14], but none have accurately captured the impact of forks on mining fairness. Consequently, the analyses conducted thus far diverged from the actual values observed in blockchain networks. Herein, we discuss each of these studies in detail.
Kanda et al. introduced the concept of effective hashrate, based on the idea that miners cannot contribute to the main chain until they receive the latest block [10]. They calculated the effective hashrate by multiplying the original hashrate by the ratio of the time required to receive a block, to the block generation interval. They asserted that the proportion of the effective hashrate is equal to the block reward rate. The problem with their calculation method is that it ignores the capability of miners to still contribute to the main chain even if they have not received the latest block. Blocks generated by miners who have not received the latest block can cause chain ties and may still be included in the main chain. Moreover, miners can increase their block reward rate by invalidating the opponent’s block during chain ties. The same problem arose in the study by Tong et al [15].
Jiang et al. analyzed mining fairness based on a concept similar to that proposed by Kanda et al [11]. They calculated the average block reception time for each miner and defined the maximum difference in the average block reception times as mining fairness. However, their concept of mining fairness evidently deviates from reality.
Xiao et al. proposed a model-based approach to analyzing mining fairness in a blockchain network for the analysis of the blockchain network connectivity [13]. The main difference between their approach and our model-based calculation is that in our method, the round start rate is taken into account. By contrast, they assumed that the round start rate equals the proportion of hashrate, which is not necessarily true.
Chen et al. examined the impact of forks on the hashrate distribution of blockchain networks from the perspective of mining fairness [14]. Their analysis had two main limitations. First, they did not consider the impact of forks on the round start rate. They assumed that the round start rate equals the proportion of the hashrate, which our analysis has demonstrated to be incorrect. Second, they considered only the block rewards for the miner who initiated the round and did not consider the rewards for subsequent blocks in the round.
Mao et al. investigated how the manner in which miners are connected affects mining fairness [16]. They conducted both theoretical analyses and simulations; however, each approach had its own limitations. With regard to their theoretical analysis, they considered only a few blocks directly connected to the genesis block. On the other hand, with regard to their simulation analysis, they ran a scenario where the ratios of the propagation time to the average block generation interval, $d/T$, were extremely high. Although this condition allows the observation of trends related to mining fairness, it does not accurately reflect the behavior of actual blockchain networks, raising concerns about the generalizability of the results.
Huang et al. compared the proof of work (PoW) and proof of stake (PoS) from the perspective of mining fairness [12]. They claimed that mining fairness is achieved in PoW blockchains in the sense that the block reward rate equals the proportion of the hashrate. However, this conclusion arose because they did not consider unintentional forks in PoW blockchains. They also analyzed the convergence rate in addition to the expected value of mining fairness. By contrast, in our study, we analyzed only the expected value of mining fairness.
Mining Fairness for Intentional Forks: Attacks that intentionally compromise mining fairness to unjustly increase block reward rates have been studied extensively [30, 4, 31, 32]. For example, Eyal et al. proposed a mining strategy called selfish mining, which increases the block reward rate by intentionally causing forks [4]. They defined the success of selfish mining as achieving a block reward rate that exceeds the proportion of the hashrate, indicating that local mining fairness becomes positive. Sun et al. introduced the Kullback–Leibler (KL) divergence between the distributions of block reward rates and proportions of hashrate as a measure of the impact of selfish mining [33]. This concept is similar to global mining fairness. However, whereas KL divergence measures the difference between distributions, what we are interested in is the distribution of the difference between block reward rates and the proportions of hashrate. In fact, the KL divergence can be zero even when mining fairness is compromised. Therefore, their definition of mining fairness has limited expressiveness. Consequently, this study does not address mining fairness based on how they defined it.
6.2 Impact of Tie-Breaking Rules on Mining Fairness
Tie-breaking rules significantly affect the probability of being included in the main chain during a fork. Our analysis indicates that the first-seen rule compromises mining fairness the most. In Section 5.2, we found that the last-generated rule leads to the highest mining fairness. By contrast, in Section 5.3, the random rule achieved the highest fairness. A more general analysis remains a subject for future research.
7 Conclusion
In this paper, we propose an efficient method for calculating mining fairness, one of the key metrics in blockchain, by approximating a complex blockchain network with a simpler network, where the number of blocks per round is at most two. We also analyzed mining fairness based on several aspects. Specifically, we examined mining fairness in a simple network composed of two miners and the impact of the hashrate distribution on mining fairness. Finally, we analyzed existing blockchain network improvement methods from the perspective of mining fairness.